OpenStudy (anonymous):

Find an exact value. sin (17pi/12)

OpenStudy (anonymous):

you can always plug it in a calculator

OpenStudy (anonymous):

idk how

OpenStudy (anonymous):

try using this websiite: http://www.wolframalpha.com/

OpenStudy (anonymous):

Type in what you wrote and it will give you the exact value.

OpenStudy (anonymous):

It will also explain some reasoning behind it.

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=sin+%2817pi%2F12%29 here i plug it in for you

OpenStudy (anonymous):

the input is wrong its not written correctly

OpenStudy (anonymous):

sin (17pi ) parenthesis around the whole thing ----- 12

OpenStudy (anonymous):

Ok then plug that in.

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

Idk you just wrote it for me. lol you tell me

OpenStudy (anonymous):

sin (17pi/12) = sin255

OpenStudy (anonymous):

i think they want her to evaluate that expression without a calculator, and to do that you would have to refer to the calculator

OpenStudy (anonymous):

i mean refer to the UNIT CIRCLE and not the calculator

OpenStudy (anonymous):

Sin ( 180 + 75) = -Sin 75

OpenStudy (anonymous):

Pretty sure that "exact value" means -not- to use a calculator but something like sin(a-b)=

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=sin+%2817pi+over+12%29 it didnt give me what i was looking for the answer choices have sqr.6 and sqr.2/4 with differnt signs as answers

OpenStudy (anonymous):

do you know what the sin and cosine of 1pi, 1/2 pi and 1/6 pi? also known as 180 degrees, 90 degrees and 30 degrees?

OpenStudy (anonymous):

Sin 75 = sin(45 + 30) = sin45 cos 30 + cos45 sin30

OpenStudy (anonymous):

no

OpenStudy (anonymous):

wat no ?

OpenStudy (anonymous):

where is that located confused?

OpenStudy (anonymous):

17pi / 12 to degrees is (17pi)/12 * 180/pi, which is 255 degrees, can yu locate 255 degrees?

OpenStudy (anonymous):

You can't just pop up a unit circle with values. You have to explain how you get those values for those certain degrees.

OpenStudy (anonymous):

fine, then explain to her

OpenStudy (anonymous):

did you see thw attacment with the question

OpenStudy (anonymous):

yea, i saw it